# On the well-posedness of SPDEs with locally Lipschitz coefficients

### v2

@@ 2. Truncation, and preliminary results
$\mathrm{P}\left\{|u_{N+1}(t\,,x)|\geq\mathrm{e}^{N}\right\}\leq\exp\left(-\frac{\mathrm{e}^{2N-4L_{b}t}}{32\mathrm{e}\left(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}+1\right)^{2}}\right),$
@@
$N\geq\frac{1}{2}\log 32+2L_{b}t+\frac{1}{2}+\log\left(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}+1\right).$
@@
Let us first consider the case that $L_{\sigma}>0$. Proposition 4 and Chebyshev’s inequality together ensure that
@@
uniformly for all real numbers $N,t>0$, $x\in\mathbb{R}$, and all real numbers $k\geq\max(2\,,L_{b}^{1/2}L_{\sigma}^{-2})$. Set $k=\sqrt{AN}$ – where the value of $A>0$ is to be determined – and $C=4(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1)$ in order to see that
@@
which has the desired outcome, when $L_{\sigma}>0$, provided additionally that $N\geq 4\log C$. The case $\sigma\in L^{\infty}(\mathbb{R})$ is proved similarly but rests on Proposition 5 instead of 4, viz.,
@@
Finally, the following real-variable lemma will be a key result for the proof of existence of the solution to (1.1).
Lemma 7.
Choose and fix two strictly positive real numbers $\alpha$ and $\gamma$ that satisfy $\gamma>\alpha^{-1}$. Then, $\sum_{N=1}^{\infty}a_{N}^{-\alpha}\exp(-N^{\gamma}/a_{N})<\infty$ for every non-decreasing sequence $0<a_{1}\leq a_{2}\leq a_{3}\leq\cdots$ of positive numbers.
@@
Choose and fix a number $\rho\in(\alpha^{-1}\!,\gamma)$ and divide the sum according to whether or not $a_{N}\leq N^{\rho}$. The lemma is true because $\sum_{a_{N}>N^{-\rho}}\leq\sum_{a_{N}>N^{\rho}}a_{N}^{-\alpha}\leq\sum_{a_{N}>N^{\rho}}N^{-\alpha\rho}$ and $\sum_{a_{N}\leq N^{-\rho}}a_{N}^{-\alpha}\exp(-N^{\gamma}/a_{N})\leq a_{1}^{-\alpha}\sum_{a_{N}\leq N^{\rho}}\exp(-N^{\gamma-\rho})$.
3. Proof of Theorem 3
@@
We will prove the theorem in the case that $L_{\sigma}>0$. Small adjustments to proof yield the result also in the remaining case that $\sigma$ is bounded.
Throughout the proof we choose and fix a non-random number $T>0$, and then strive to prove that (1.1) has a mild solution up to time $T$. Since $T$ is arbitrary, this proves the theorem. With this number in mind, we adjust (2.7) slightly and define
$\mathcal{N}_{k,\beta,T}(Z)=\sup_{t\in(0,T)}\sup_{x\in\mathbb{R}}\mathrm{e}^{-\beta t}\|Z(t\,,x)\|_{k},$
for every $\beta>0$, $k\geq 1$, and all space-time random fields $Z$.
Note that $N\mapsto\text{\rm Lip}_{N}(b)$ and $N\mapsto\text{\rm Lip}_{N}(\sigma)$ are nondecreasing and $b$ and $\sigma$ are globally Lipschitz when $\lim_{N\to\infty}\text{\rm Lip}_{N}(b)<\infty$ and $\lim_{N\to\infty}\text{\rm Lip}_{N}(\sigma)<\infty$. Therefore, we assume without loss of generality that
$\lim_{N\to\infty}L_{N,b}=\infty\quad\text{and}\quad\lim_{N\to\infty}L_{N,\sigma}=\infty,$ (3.1)
where $L_{N,b}$ and $L_{N,\sigma}$ were defined in (2.1).
Thanks to (2.4) we can write
@@
$\leq\left\|b_{N}(s\,,u_{N+1}(s\,,y))-b_{N}(s\,,u_{N}(s\,,y))\right\|_{k}$
$\leq L_{N,b}\left\|u_{N+1}(s\,,y)-u_{N}(s\,,y)\right\|_{k}{\color[rgb]{0,0,0}\leq L_{N,b}\mathrm{e}^{\beta s}\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N}),}$
for all $\beta,N>0$, $s\in(0\,,T)$, and $y\in\mathbb{R}$. On the other hand,
@@
We have used the following variation of the Cauchy–Schwarz inequality in the last line: $\|X\mathbf{1}_{F}\|_{k}\leq\|X\|_{2k}[\mathrm{P}(F)]^{1/(2k)}$ for all $X\in L^{k}(\Omega)$ and all events $F\subset\Omega$. Set $C=4(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1).$ Thanks to Proposition 4,
@@
$\leq L_{b}\left[\|u_{N+1}(s\,,y)\|_{2k}+\|u_{N}(s\,,y)\|_{2k}\right]\leq{\color[rgb]{0,0,0}2CL_{b}}\mathrm{e}^{512L_{\sigma}^{4}k^{2}s},$
@@
$\qquad\leq{\color[rgb]{0,0,0}2CL_{b}}\mathrm{e}^{512L_{\sigma}^{4}k^{2}s}\left[\mathrm{P}\left\{|u_{N+1}(s\,,y)|\geq\mathrm{e}^{N}\right\}\right]^{1/(2k)},$
valid uniformly for all $N,s>0$, $y\in\mathbb{R}$, and $k\geq\max(1\,,\frac{1}{2}L_{b}^{1/2}L_{\sigma}^{-2})$. Therefore, Proposition 6 yields
@@
$\qquad\leq 2L_{b}C\mathrm{e}^{512L_{\sigma}^{4}k^{2}s}\exp\left(-\frac{N^{3/2}}{2k(512s)^{1/2}L_{\sigma}^{2}}\right),$
valid uniformly for all $s\in(0\,,T)$, $y\in\mathbb{R}$, $N\geq 4\log C\vee 256T\max\left(4L_{\sigma}^{4},L_{b}\right)$, and $k\geq\max(1\,,\frac{1}{2}L_{b}^{1/2}L_{\sigma}^{-2})$. Thus, we find that
@@
$\quad+2L_{b}C\int_{(0,t)\times\mathbb{R}}\mathrm{e}^{512L_{\sigma}^{4}k^{2}s}\exp\left(-\frac{N^{3/2}}{2k(512s)^{1/2}L_{\sigma}^{2}}\right)p_{t-s}(y-x)\,\mathrm{d}s\,\mathrm{d}y$ (3.3)
$\leq\frac{L_{N,b}\mathrm{e}^{\beta t}}{\beta}\,\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N})+\frac{L_{b}C}{256L_{\sigma}^{4}k^{2}}\,\exp\left(512L_{\sigma}^{4}k^{2}t-\frac{N^{3/2}}{2k(512t)^{1/2}L_{\sigma}^{2}}\right),$
uniformly for all $t\in(0\,,T)$ and $x\in\mathbb{R}$, provided that $N,k\geq c$, for a complicated looking but otherwise unimportant number $c=c(\|u_{0}\|_{L^{\infty}(\mathbb{R})}\,,L_{\sigma}\,,T\,,L_{b})>0$.
@@
$\leq L_{N,\sigma}\mathrm{e}^{\beta s}\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N})\quad\text{ for all $\beta,N>0$, $s\in(0\,,T)$, and $y\in\mathbb{R}$.}$
@@
$\hskip 7.22743pt\leq\left[\left\|\sigma_{N+1}(s\,,u_{N+1}(s\,,y))\right\|_{2k}+\left\|\sigma_{N}(s\,,u_{N}(s\,,y))\right\|_{2k}\right]\left[1-\mathrm{P}\left(G_{N+1}(s\,,y)\right)\right]^{1/(2k)}.$
Thanks to Proposition 4,
@@
uniformly for all $N,s>0$, $y\in\mathbb{R}$, and $k\geq\max(1\,,\frac{1}{2}L_{b}^{1/2}L_{\sigma}^{-2})$. Therefore, Proposition 6 yields
@@
$\hskip 122.85876pt\leq 2L_{\sigma}C\mathrm{e}^{512L_{\sigma}^{4}k^{2}s}\exp\left(-\frac{N^{3/2}}{2k(512s)^{1/2}L_{\sigma}^{2}}\right),$
uniformly for all $s\in(0\,,T)$, $y\in\mathbb{R}$, $N\geq 4\log C\vee 256T\max\left(4L_{\sigma}^{4},L_{b}\right)$, and $k\geq\max(1\,,\frac{1}{2}L_{b}^{1/2}L_{\sigma}^{-2})$. Thus, we find that
@@
$\quad+32kL_{b}^{2}C^{2}\int_{(0,t)\times\mathbb{R}}\mathrm{e}^{1024L_{\sigma}^{4}k^{2}s}\exp\left(-\frac{N^{3/2}}{k(512s)^{1/2}L_{\sigma}^{2}}\right)[p_{t-s}(y-x)]^{2}\,\mathrm{d}s\,\mathrm{d}y$
$\leq\frac{4kL_{N,\sigma}^{2}\mathrm{e}^{2\beta t}}{\sqrt{\beta}}\left[\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N})\right]^{2}{\color[rgb]{0,0,0}+32kL_{\sigma}^{2}C^{2}\mathrm{e}^{1024L_{\sigma}^{4}k^{2}t}\exp\left(-\frac{N^{3/2}}{k(512t)^{1/2}L_{\sigma}^{2}}\right)\sqrt{t},}$
for every choice of $\beta>0$, $t\in(0\,,T)$, and $x\in\mathbb{R}$, provided that $N,k\geq c$. Combine this last inequality with (3.3) and (3.2) in order to see that
@@
$\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{k}\leq\mathrm{e}^{\beta t}\left[\frac{L_{N,b}}{\beta}+\frac{2\sqrt{k}L_{N,\sigma}}{\beta^{1/4}}\right]\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N})$
$\qquad+\frac{L_{b}C}{256L_{\sigma}^{4}k^{2}}\,\exp\left(512L_{\sigma}^{4}k^{2}t-\frac{N^{3/2}}{2k(512t)^{1/2}L_{\sigma}^{2}}\right)$
$\qquad+\sqrt{32k}L_{\sigma}C\mathrm{e}^{512L_{\sigma}^{4}k^{2}t}\exp\left(-\frac{N^{3/2}}{2k(512t)^{1/2}L_{\sigma}^{2}}\right)t^{1/4}.$
as long as $N,k\geq c$. Set $\beta=512L_{N,\sigma}^{4}k^{6}$, divide both sides by $\exp(\beta t)$ and optimize over $(t\,,x)$ in order to see that
@@
$\mathcal{N}_{k,512L_{N,\sigma}^{4}k^{6},T}(u_{N+1}-u_{N})\leq\left[\frac{L_{N,b}}{4^{4}L_{N,\sigma}^{4}k^{6}}+\frac{1}{4k^{11/2}}\right]\mathcal{N}_{k,512L_{\sigma}^{4}k^{6},T}(u_{N+1}-u_{N})$
$\qquad\qquad+\frac{L_{b}C}{256L_{\sigma}^{4}k^{2}}\,\exp\left(512t(L_{\sigma}^{4}k^{2}-L_{N,\sigma}^{4}k^{6})-\frac{N^{3/2}}{2k(512T)^{1/2}L_{\sigma}^{2}}\right)$ (3.4)
$\qquad\qquad+\sqrt{32k}t^{1/4}L_{\sigma}C\mathrm{e}^{512t(L_{\sigma}^{4}k^{2}-L_{N,\sigma}^{4}k^{6})}\exp\left(-\frac{N^{3/2}}{2k(512T)^{1/2}L_{\sigma}^{2}}\right).$
Set $k=\sqrt{L_{N,b}}$ and take $N$ large enough such that $\alpha_{0}:=(4^{4}L_{N,\sigma}^{4}L_{N,b}^{2})^{-1}+(4L_{N,b}^{11/4})^{-1}<1$ and $L_{\sigma}^{4}<L_{N,\sigma}^{4}$. In particular, if $k^{4}\geq 2$, we observe that $\exp\{512t(L_{\sigma}^{4}k^{2}-L_{N,\sigma}^{4}k^{6})\}\leq\exp\{-512tL_{N,\sigma}^{4}k^{6}(1-\frac{1}{k^{4}})\}\leq\exp\{-256tk^{6}L_{N,\sigma}^{4}\}$. Since $\exp(-x)\leq x^{-1/4}$ when $x>0$, we find that
$\adjustlimits{\sup}_{t\in(0,T)}{\sup}_{x\in\mathbb{R}}\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{\sqrt{L_{N,b}}}$
$\hskip 36.135pt\leq\frac{1}{(1-\alpha_{0})}\frac{L_{b}C}{256L_{\sigma}^{4}L_{N,b}}\,\exp\left(-\frac{N^{3/2}}{2\sqrt{L_{N,b}}(512T)^{1/2}L_{\sigma}^{2}}\right)$
$\hskip 50.58878pt+\frac{1}{(1-\alpha_{0})}\frac{2C}{L_{N,b}}\exp\left(-\frac{N^{3/2}}{2\sqrt{L_{N,b}}(512T)^{1/2}L_{\sigma}^{2}}\right),$
provided only that $N,k\geq c$ [^1] It is possible to use a priori estimates to show that the above is the only possible subsolution to (3.4); in other words, that $\mathcal{N}_{k,512L_{N,\sigma}^{4}k^{2},T}(u_{N+1}-u_{N})$ is not infinite. The method uses Picard iteration and follows along similar lines that were developed in Foondun and Khoshnevisan [FK1]. We skip the several lines of additional technical details. Thanks to (3.1), the above bounds $\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{2}$ for all large $N$, and we find that
$\adjustlimits{\sup}_{t\in(0,T)}{\sup}_{x\in\mathbb{R}}\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{2}\lesssim L_{N,b}^{-1}\exp\left(-\frac{N^{3/2}}{2(512T)^{1/2}L_{\sigma}^{2}\sqrt{L_{N,b}}}\right),$
uniformly for all $N\geq c^{\prime}=c^{\prime}(\|u_{0}\|_{L^{\infty}(\mathbb{R})}\,,L_{\sigma}\,,L_{b}\,,T)>0$ for a complicated looking but otherwise unimportant number $c^{\prime}$. We appeal to Lemma 7 [with $a_{N}=2(512T)^{1/2}L_{\sigma}^{2}\sqrt{L_{N,b}}$, $\gamma=3/2$ and $\alpha=2$] to see that
$\sum_{\begin{subarray}{c}N\in\mathbb{N}\cap[c^{\prime},\infty)\end{subarray}}\adjustlimits{\sup}_{t\in(0,T)}{\sup}_{x\in\mathbb{R}}\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{2}<\infty.$
Thus, we may deduce from (2.3) and (3.1) that the above continues to hold if we replace the sum with the sum over all $N\in\mathbb{N}$. This proves that
$u(t\,,x)=\lim_{N\to\infty}u_{N}(t\,,x)\quad\text{exists in $L^{2}(\Omega)$},$ (3.5)

### v3

@@ 2. Truncation, and preliminary results
$\mathrm{P}\left\{|u_{N+1}(t\,,x)|\geq\mathrm{e}^{N}\right\}\leq\exp\left(-\frac{\mathrm{e}^{2N-4L_{b}t}}{32\mathrm{e}\left(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}+1\right)^{2}}\right),$ (2.11)
@@
$N\geq\tfrac{1}{2}\log 32+2L_{b}t+\tfrac{1}{2}+\log\left(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}+1\right).$
@@
Let us first consider the case that $L_{\sigma}>0$. Proposition 2.1 and Chebyshev’s inequality together ensure that
@@
uniformly for all real numbers $N,t>0$, $x\in\mathbb{R}$, and all real $k\geq\max(2\,,L_{b}^{1/2}L_{\sigma}^{-2})$. Set $k=\sqrt{AN}$ – where the value of $A>0$ is to be determined – and $C=4(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1)$ in order to see that
@@
which has the desired outcome, when $L_{\sigma}>0$, provided additionally that $N\geq 4\log C$. The case $\sigma\in L^{\infty}(\mathbb{R})$ is proved similarly but rests on Proposition 2.3 instead of 2.1, viz.,
@@
Finally, the following real-variable lemma will be helpful to us.
Lemma 2.5.
[…]
$\sup_{t\in(0,T]}\left[\mathrm{e}^{-\beta t}f(t)\right]\leq\mathrm{e}^{-\beta T}g(T)\qquad\forall T\in(0\,,T_{0}),$
then $\sup_{(0,T]}f\leq g(T)$ for every $T\in(0\,,T_{0})$.
@@
Since $\mathrm{e}^{-\beta T}f(T)\leq\mathrm{e}^{-\beta T}g(T)$, we cancel the exponentials to deduce the result from the monotonicity of $g$. ∎
3. Proof of Theorem 1.4
@@
We first prove the theorem in the case that $L_{\sigma}>0$. Until we mention to the contrary, we therefore assume tacitly that $L_{\sigma}>0$. Thus, we also assume that (1.6) holds.
Let us choose and fix an arbitrary number $T>0$. Our primary goal is to prove that
$\sum_{N=1}^{\infty}\adjustlimits{\sup}_{t\in[0,T]}{\sup}_{x\in\mathbb{R}}\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{k}<\infty\qquad\forall k\geq 1.$ (3.1)
Because $T>0$ can be as large as needed, (3.1) implies that the random variable $u(t\,,x)=\lim_{N\to\infty}u_{N}(t\,,x)$ exists in $L^{2}(\Omega)$, say, pointwise in (t,x)(t\,,x). From there it is not difficult to prove that the thus-defined random field $u$ is a mild solution to (1.1).
With the preceding paragraph in mind, we now adjust (2.5) slightly and define the following local-in-time norms:
$\mathcal{N}_{k,\beta,T}(Z)=\adjustlimits{\sup}_{t\in(0,T]}{\sup}_{x\in\mathbb{R}}\mathrm{e}^{-\beta t}\|Z(t\,,x)\|_{k},$ (3.2)
defined for every $\beta,T>0$, $k\geq 1$, and all space-time random fields $Z$.
Recall (1.5) and note that $N\mapsto\text{\rm Lip}_{N}(b)$ and $N\mapsto\text{\rm Lip}_{N}(\sigma)$ are nondecreasing and $b$ and $\sigma$ are globally Lipschitz respectively when $\lim_{N\to\infty}\text{\rm Lip}_{N}(b)<\infty$ and $\lim_{N\to\infty}\text{\rm Lip}_{N}(\sigma)<\infty$. Therefore, Condition (1.6) ensures that we can assume without any loss in generality that
$\lim_{N\to\infty}L_{N,\sigma}=\infty,$ (3.3)
for $b$ and $\sigma$ will be globally Lipschitz otherwise, in which case there is nothing to prove. From now on, condition (3.3) is assumed to hold.
Thanks to (2.2) we can write, for every $k\geq 1$ [not necessarily an integer], $t>0$, and $x\in\mathbb{R}$,
@@
$\hskip 72.26999pt\leq\left\|b_{N}(s\,,u_{N+1}(s\,,y))-b_{N}(s\,,u_{N}(s\,,y))\right\|_{k}$
$\hskip 72.26999pt\leq L_{N,b}\left\|u_{N+1}(s\,,y)-u_{N}(s\,,y)\right\|_{k}\leq L_{N,b}\mathrm{e}^{\beta s}\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N}),$
for all $k\geq 1$, $\beta,N>0$, $s\in(0\,,T]$, and $y\in\mathbb{R}$. On the other hand,
@@
We have used the following variation of the Cauchy–Schwarz inequality in the last line: $\|X\mathbf{1}_{F}\|_{k}\leq\|X\|_{2k}[\mathrm{P}(F)]^{1/(2k)}$ for all $X\in L^{k}(\Omega)$ and all events $F\subset\Omega$. Set $C=4(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1).$ Thanks to Proposition 2.1,
@@
$\hskip 72.26999pt\leq L_{b}\left[\|u_{N+1}(s\,,y)\|_{2k}+\|u_{N}(s\,,y)\|_{2k}\right]\leq 2CL_{b}\mathrm{e}^{512L_{\sigma}^{4}k^{2}s},$
@@
$\hskip 72.26999pt\leq 2CL_{b}\mathrm{e}^{512L_{\sigma}^{4}k^{2}s}\left[\mathrm{P}\left\{|u_{N+1}(s\,,y)|\geq\mathrm{e}^{N}\right\}\right]^{1/(2k)},$
valid uniformly for all $N,s>0$, $y\in\mathbb{R}$, and $k\geq 1$. Therefore, Proposition 2.4 yields
@@
$\hskip 180.67499pt\leq 2L_{b}C\mathrm{e}^{512L_{\sigma}^{4}k^{2}s}\exp\left(-\frac{N^{3/2}}{128kL_{\sigma}^{2}\sqrt{s}}\right),$
valid uniformly for all $s\in(0\,,T]$, $y\in\mathbb{R}$, $N\geq 4\log C\vee 256T\max(4L_{\sigma}^{4},L_{b})$, and $k\geq 1$. Thus, we find using (3.2) that
@@
$\quad+2L_{b}C\int_{(0,t)\times\mathbb{R}}\mathrm{e}^{512L_{\sigma}^{4}k^{2}s}\exp\left(-\frac{N^{3/2}}{128kL_{\sigma}^{2}\sqrt{s}}\right)p_{t-s}(y-x)\,\mathrm{d}s\,\mathrm{d}y$ (3.5)
$\leq\frac{L_{N,b}\mathrm{e}^{\beta t}}{\beta}\,\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N})+\frac{L_{b}C}{128L_{\sigma}^{4}k^{2}}\,\exp\left(512L_{\sigma}^{4}k^{2}t-\frac{N^{3/2}}{128kL_{\sigma}^{2}\sqrt{t}}\right),$
uniformly for all $t\in(0\,,T)$ and $x\in\mathbb{R}$, provided that $N,k\geq c$, for a complicated looking but otherwise unimportant number
$c=c(\|u_{0}\|_{L^{\infty}(\mathbb{R})}\,,L_{\sigma}\,,T\,,L_{b})>1.$ (3.6)
For later purposes, we pause to mention that the number $c$, while fixed, can be chosen to be as large as we wish. Owing to Condition (1.6), we therefore select $c=c(\|u_{0}\|_{L^{\infty}(\mathbb{R})}\,,L_{\sigma}\,,T\,,L_{b})$ large enough to additionally ensure that
$c>\sup_{N\geq N_{0}}\frac{\sqrt{L_{N,b}}}{L_{N,\sigma}^{2}},\quad\text{where}\quad N_{0}=\inf\left\{N>1:\ L_{N,\sigma}\geq 1\right\}.$ (3.7)
The number $N_{0}$ is well defined and finite thanks to (3.3).
@@
$\leq L_{N,\sigma}\mathrm{e}^{\beta s}\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N})\qquad\forall\beta,N>0,\ s\in(0\,,T],\ y\in\mathbb{R}.$
@@
$\quad\leq\left[\left\|\sigma_{N+1}(s\,,u_{N+1}(s\,,y))\right\|_{2k}+\left\|\sigma_{N}(s\,,u_{N}(s\,,y))\right\|_{2k}\right]\left[1-\mathrm{P}\left(G_{N+1}(s\,,y)\right)\right]^{1/(2k)}.$
Thanks to Proposition 2.1,
@@
uniformly for all $N,s>0$, $y\in\mathbb{R}$, and $k\geq 1$. Therefore, Proposition 2.4 yields
@@
$\hskip 122.85876pt\leq 2L_{\sigma}C\mathrm{e}^{512L_{\sigma}^{4}k^{2}s}\exp\left(-\frac{N^{3/2}}{128kL_{\sigma}^{2}\sqrt{s}}\right),$
uniformly for all $s\in(0\,,T]$, $y\in\mathbb{R}$, $k\geq 1$, and
$N\geq c_{T}:=4\log C\vee 256T\max\left(4L_{\sigma}^{4},L_{b}\right).$ (3.8)
Thus, we find that
@@
$\quad+32kL_{b}^{2}C^{2}\int_{(0,t)\times\mathbb{R}}\mathrm{e}^{1024L_{\sigma}^{4}k^{2}s}\exp\left(-\frac{N^{3/2}}{128kL_{\sigma}^{2}\sqrt{s}}\right)[p_{t-s}(y-x)]^{2}\,\mathrm{d}s\,\mathrm{d}y$
@@
$\leq\frac{4kL_{N,\sigma}^{2}\mathrm{e}^{2\beta t}}{\sqrt{\beta}}\left[\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N})\right]^{2}+128kL_{\sigma}^{2}C^{2}\mathrm{e}^{1024L_{\sigma}^{4}k^{2}t}\exp\left(-\frac{N^{3/2}}{128kL_{\sigma}^{2}\sqrt{t}}\right)\sqrt{t},$
for every choice of $\beta>0$, $t\in(0\,,T)$, and $x\in\mathbb{R}$, provided that $N\geq\max(c\,,c_{T})$ and $k\geq c$ – in case it helps we recall that $c>1$ and $c_{T}>0$ were defined respectively in (3.6) and (3.8). Combine this last inequality with (3.5) and (3.4) in order to see that
$\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{k}\leq\mathrm{e}^{\beta t}\left[L_{N,b}\beta^{-1}+2\sqrt{k}L_{N,\sigma}\beta^{-1/4}\right]\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N})$
$\hskip 72.26999pt+C\left[\frac{L_{b}}{128L_{\sigma}^{4}k^{2}}+\sqrt{128k}t^{1/4}L_{\sigma}\right]\exp\left(512L_{\sigma}^{4}k^{2}t-\frac{N^{3/2}}{128kL_{\sigma}^{2}\sqrt{t}}\right),$
as long as $N\geq\max(c\,,c_{T})$ and $k\geq c$.
Next, we make particular choices of $k\geq c$ and $\beta>0$ as follows:
$k=c\quad\text{and}\quad\beta=16A_{0}^{4}k^{2}L_{N,\sigma}^{4},$ (3.9)
where
$A_{0}=\max\left(\sqrt{8}L_{\sigma}^{4}\,,4\right).$ (3.10)
We pause to emphasize that $\beta$ depends on (N,T)(N\,,T). In this way we find that
@@
$\sup_{x\in\mathbb{R}}\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{c}\leq\mathrm{e}^{\beta t}\left[\frac{L_{N,b}}{16A_{0}^{4}c^{2}L_{N,\sigma}^{4}}+\frac{1}{A_{0}}\right]\mathcal{N}_{c,\beta,T}(u_{N+1}-u_{N})$
$\hskip 72.26999pt+C\left[\frac{L_{b}}{128L_{\sigma}^{4}c^{2}}+\sqrt{128c}t^{1/4}L_{\sigma}\right]\exp\left(512L_{\sigma}^{4}c^{2}t-\frac{N^{3/2}}{128cL_{\sigma}^{2}\sqrt{t}}\right),$
uniformly for all $T>0$, $t\in(0\,,T]$, and $N\geq\max(N_{0}\,,c\,,c_{T})$; see (3.7). Because $A_{0}\geq 4$ – see (3.10) – an application of (3.7) yields the following inequality,
$\sup_{x\in\mathbb{R}}\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{c}\leq\mathrm{e}^{\beta t}\left[\frac{1}{16A_{0}^{4}}+\frac{1}{A_{0}}\right]\mathcal{N}_{c,\beta,T}(u_{N+1}-u_{N})$
$\hskip 72.26999pt+C\left[\frac{L_{b}}{128L_{\sigma}^{4}c^{2}}+\sqrt{128c}t^{1/4}L_{\sigma}\right]\exp\left(512L_{\sigma}^{4}c^{2}t-\frac{N^{3/2}}{128cL_{\sigma}^{2}\sqrt{t}}\right)$
$\leq\mathrm{e}^{\beta t}\left[\frac{1}{4096}+\frac{1}{4}\right]\mathcal{N}_{c,\beta,T}(u_{N+1}-u_{N})$
$\hskip 72.26999pt+C\left[\frac{L_{b}}{128L_{\sigma}^{4}c^{2}}+\sqrt{128c}t^{1/4}L_{\sigma}\right]\exp\left(512L_{\sigma}^{4}c^{2}t-\frac{N^{3/2}}{128cL_{\sigma}^{2}\sqrt{t}}\right),$
valid uniformly for all $T>0$, $t\in(0\,,T]$, and $N\geq\max(N_{0}\,,c\,,c_{T})$. Since $\frac{1}{4096}+\frac{1}{4}<\frac{1}{2}$, we may divide both sides of the preceding by $\exp(\beta t)$ and optimize over $t\in(0\,,T]$ in order to find that
$\mathcal{N}_{c,\beta,T}(u_{N+1}-u_{N})$ (3.11)
$\leq 2C\left[\frac{L_{b}}{128L_{\sigma}^{4}c^{2}}+\sqrt{128c}T^{1/4}L_{\sigma}\right]\sup_{t\in(0,T]}\exp\left(-\left(\beta-512L_{\sigma}^{4}c^{2}\right)t-\frac{N^{3/2}}{128cL_{\sigma}^{2}\sqrt{t}}\right),$
uniformly for all $T>0$, and $N\geq\max(N_{0}\,,c\,,c_{T})$. Thanks to (3.7), (3.9), and (3.10),
$\beta-512L_{\sigma}^{4}c^{2}=c^{2}\left(16A_{0}^{4}L_{N,\sigma}^{4}-512L_{\sigma}^{4}\right)\geq c^{2}\left(16A_{0}^{4}-512L_{\sigma}^{4}\right)>0,$
uniformly for all $N\geq N_{0}$. Now, let us consider the following function that appears in the exponent on the right-hand side of (3.11):
$\psi(t)=\left(\beta-512L_{\sigma}^{4}c^{2}\right)t+\frac{N^{3/2}}{128kL_{\sigma}^{2}\sqrt{t}}\qquad(t>0).$
Because
$\psi^{\prime}(t)\leq\beta-\frac{(N/T)^{3/2}}{256cL_{\sigma}^{2}}=16A_{0}^{4}L_{N,\sigma}^{4}c^{2}-\frac{(N/T)^{3/2}}{256cL_{\sigma}^{2}}\quad\forall 0<t\leq T,$
it follows from this that $\psi^{\prime}<0$ everywhere on (0,T](0\,,T] provided that
$\frac{N}{L_{N,\sigma}^{8/3}}>4096^{2/3}A_{0}^{8/3}c^{2}L_{\sigma}^{4/3}T=:N_{T};$ (3.12)
the left-hand side being well defined for example when $N_{T}>N_{0}$. Condition (1.6) ensures that the left-hand side tends to infinity as $N\to\infty$. Therefore, (3.12) holds for every $N>\max(N_{0}\,,N_{T})$. It follows that
$\inf_{t\in(0,T]}\psi(t)=\psi(T)\quad\forall N>\max(N_{0}\,,N_{T}),\ T>0,$
whence it follows from (3.11) that, for every $T_{0}>0$ and $N\geq\max(N_{0}\,,N_{T_{0}}\,,c_{0}\,,c_{T_{0}})$,
$\mathcal{N}_{c,\beta,T}(u_{N+1}-u_{N})$
$\leq 2C\left[\frac{L_{b}}{128L_{\sigma}^{4}c^{2}}+\sqrt{128c}T^{1/4}L_{\sigma}\right]\exp\left(-\left(\beta-512L_{\sigma}^{4}c^{2}\right)T-\frac{N^{3/2}}{128cL_{\sigma}^{2}\sqrt{T}}\right),$
uniformly for every $T\in(0\,,T_{0})$. In order to ensure the uniformity statement of $T$, we have also used the fact that $T\mapsto c_{T}$ is increasing; see (3.8). We now apply Lemma 2.5 in order to deduce from the above and (3.2) that, for every $T>0$ fixed,
$\limsup_{N\to\infty}N^{-3/2}\log\adjustlimits{\sup}_{t\in(0,T]}{\sup}_{x\in\mathbb{R}}\|u_{N+1}(t\,,x)-u_{N}(t)\|_{k}<0\qquad\forall k\geq 1.$ (3.13)
When $k\in[1\,,c]$ this follows from the preceding. For general $k$ it follows from a relabeling $k\leftrightarrow c$], and an appeal to the fact that $c$ can be as large as we wish – see the comments that follow (3.6). This proves our original goal (3.1). The remainder of the argument is technically simpler.
The preceding proves that
$u(t\,,x)=\lim_{N\to\infty}u_{N}(t\,,x)\quad\text{exists in $L^{k}(\Omega)$}\qquad\forall k\geq 1,$ (3.14)